9 quarts of a sugar solution was mixed 6 quarts of a 90% sugar solution to make an 84% sugar solution. What is the percent concentration of the first solution?
Answer by ankor@dixie-net.com(22740)

9 quarts of a sugar solution was mixed 6 quarts of a 90% sugar solution
to make an 84% sugar solution.
What is the percent concentration of the first solution?
:
Let x = the decimal equiv of the required percentage of the 1st solution
The total amt: 9 + 6 = 15 qts
:
9x + .90(6) = .84(15)
9x + 5.4 = 12.6
9x = 12.6-5.4
9x = 7.2
x = 7.2/9
x = .8 which is an 80% solution
: